# Audio dequantization using instantaneous frequency

### v1

@@ Abstract
We present a dequantization method that employs a phase-aware regularizer, originally successfully applied in an audio inpainting problem. The method maintains the temporal continuity of sinusoidal components in the audio signal time-frequency representation and avoids the energy loss artifacts commonly encountered with $\ell_{1}$-based regularization approaches. The proposed method is called the Phase-Aware Audio Dequantizer (PHADQ). The methods are evaluated using the objective metric SDR and PEMO-Q ODG.
@@ 1 Introduction
Quantization is a nonlinear distortion that arises from rounding during the A/D conversion or bit-depth reduction. This distortion appears as the quantization noise, which, at low bit depths (less than 16 bits per sample), results in unpleasant auditory artifacts [1]. Dequantization is the attempt to recover a signal close to its original form, based on the observed quantized samples. The problem is ill-conditioned; therefore, reconstruction involves a form of regularization. For instance, is the application of sparsity [2] is one potential approach. Autoregressive modeling is another modeling approach to dequantization [3, 4, 5].
In this paper, we present a dequantization method that employs a phase-aware regularizer introduced in [6], originally successfully applied in an audio inpainting problem. The method maintains the temporal continuity of sinusoidal components in the audio signal time–frequency representation and avoids the energy loss artifacts commonly encountered with $\ell_{1}$-based regularization approaches [6, 7].
@@ 2 Review of PHAIN (Phase-aware audio inpainter)
Problem (1) contains a penalty $\|DR_{\omega_{\mathbf{s}}}G_{g}{\mathbf{x}}\|_{1}$ weighted by a positive $\lambda$, and an indicator function $\imath_{\mathit{\Gamma}}({\mathbf{x}})$ taking value $\infty$ when ${\mathbf{x}}\notin\Gamma$, thus enforcing the estimation to retain the reliable parts of the signal corresponding to the set of feasible signals $\Gamma$. The linear operators in the penalty function are:
  • The discrete Gabor transform (DGT) $G$ with a window $g$.
@@
    where $a$ denotes the hop-size of the Gabor transform, and $n,m$ represent the time and frequency indices, respectively.
@@
The convex $\ell_{1}$-norm denoted $\|\cdot\|_{1}$ quantifies time-directional variation of the phase-corrected time-frequency representation of $\mathbf{x}$}. Therefore, when used as a penalty, it promotes sinusoidal components and, conversely, attenuating components whose phase cannot be adequately approximated as a linear function, such as noise or transients. As a consequence, PHAIN to a great extent solves the problem encountered in plain sparsity-based methods which is loosing signal energy inside the gap [7]. When $\omega_{\mathbf{s}}$ is fixed, the problem (1) is convex and solvable via the Chambolle–Pock algorithm [11, 12].
Based on the regime of treating the instantaneous frequency during the reconstruction, PHAIN can further be classified as follows:
  • B-PHAIN, which uses the instantaneous frequency $\omega_{\mathbf{y}}$ of the degraded signal including its unreliable parts.
  • U-PHAIN, which updates the instantaneous frequency according to the current estimate of $\hat{{\mathbf{x}}}$.
@@ 3 Method
The call the proposed method simply the Phase-Aware Audio Dequantizer (PHADQ). In the method, we employ the same penalty as in PHAIN, but for a different restoration task.
Let ${\mathbf{y}}^{\text{q}}\in\mathbb{R}^{L}$ be a quantized observation of an original audio signal with a quantization step $\Delta$. When employing uniform quantization, the quantization step size is given by $\Delta=2^{-w+1}$, where $w$ denotes the word length in bits per sample (bps). We recover the signal by solving the same convex problem as in (1), except that no reliable parts of the signal are available, and therefore the set of feasible signals is now characterized by:
@@
By choosing to update or not to update the instantaneous frequency estimate, we distinguish between the B-PHADQ and U-PHADQ, respectively.
@@ 3.1 B-PHADQ
The consistent and inconsistent problems (problems (1) and (5), respectively) are solved by Algorithm 1, which constitutes a specific instance of the Chambolle–Pock algorithm, wherein ${\mathbf{x}}\in\mathbb{R}^{L}$, ${\mathbf{p}}\in\mathbb{R}^{L}$ is a primal variable, ${\mathbf{q}}\in\mathbb{C}^{P}$ is a dual variable and ${\mathbf{u}}\in\mathbb{R}^{L}$ is a temporary variable. Positive scalars $\tau$, $\sigma$ are the step sizes and $\rho$ is the relaxation factor [12]. The scalar $\lambda>0$ is the weighting factor for the penalty function, having also an effect on the reconstruction quality and the convergence speed. Before applying the algorithm, we need to compute the instantaneous frequency $\omega_{\mathbf{s}}$, which in the case of B-PHADQ corresponds to the quantized observation, following [6]:
$\omega_{\mathbf{s}}[m,n]=\omega_{\mathbf{y}^{\text{q}}}[m,n]=-\text{Im}\Bigg[\cfrac{(G_{g^{\prime}}{\mathbf{y}^{\text{q}}})[m,n]}{(G_{g}{\mathbf{y}^{\text{q}}})[m,n]}\Bigg],$ (6)
where $G$ denotes the Gabor transform, $g^{\prime}$ is the time derivative of the window function, and $\text{Im}[\cdot]$ denotes the imaginary part.
@@
The algorithm utilizes proximal operators of the functions present in the minimization, which in the case of the $\ell_{1}$ norm correspond to the soft-thresholding, in the case of the consistent indicator function it is the projection onto the set $\Gamma$. For the inconsistent case it is given by
@@
The asterisk denotes the adjoint of a linear operator. In the algorithm 1, the term ${\mathbf{p}}^{(i+1)}$ is defined differently for the consistent and inconsistent variant. Finally, $\text{clip}_{\lambda}$ is defined as $\text{clip}_{\lambda}=\text{Id}-\text{soft}_{\lambda}$.
@@ 3.2 U-PHADQ
As in U-PHAIN [6], an alternative approach to problems (1) and (5) is to regularly update the instantaneous frequency employed in the algorithm 1. This is done after a prescribed number of iterations. The new instantaneous frequency is calculated for an estimate $\mathbf{p}$}, rather than $\mathbf{y}^{\text{q}}$, according to the current reconstruction. U-PHADQ therefore contains Algorithm 1 within an additional loop, in the same way as in U-PHAIN [6].
@@ 4 Experiment and Results
Audio has been peak-normalized before any processing to make the most of the available dynamic range. The audio was quantized to seven different word lengths, ranging from 2 to 8 bps using the mid-riser uniform quantization. In the experiment, the discrete Gabor transform was used with a Hann window of 8,192 samples in length, 75% overlap and 16,384 frequency channels. All computation was performed in MATLAB R2025a.
The parameters of Algorithm 1 were manually tuned. We used $\tau=1,\ \sigma=1$ and $\rho=1/3$. For better results, the balancing parameter $\lambda$ of the task needs to be set depending on the word length; we set it as shown in Table 1.
Table 1: Set of $\lambda$ according to word length
@@
$\lambda$ $10^{-1}$ $10^{-1}$ $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-4}$ $10^{-4}$
In the experiments, the performance of the methods was evaluated using the objective metric SDR, which expresses the signal-to-distortion ratio and measures the physical similarity of waveforms. We also employed the PEMO-Q objective difference grade (ODG) [13], which considers the perceptual characteristics of the audio signals. The ODG is measured on a scale from 4-4 to 0 (worst to best).
@@
When comparing the U-PHADQ and B-PHADQ variants, it was found that updating the instantaneous frequency within the course of iterations of U-PHADQ does not lead to better results in terms of the SDR. For this reason, in the experiments we focus only on the B-PHADQ variant. In addition, to see the maximum achievable performance of PHADQ, there is the B-PHADQ (oracle) variant, based on the instantaneous frequency of the original signal, which cannot be accessed in practical scenarios. For the oracle variant the consistent B-PHADQ was used. An example showing the evolution of SDR across iterations of PHADQ variants is presented in Fig. 1.
The results of the PHADQ method for both datasets show the best achievable SDR and ODG from a maximum of 200 iterations. The CP results were obtained after 500 iterations of the algorithm, achieving optimal ODG values [2].
@@
4.1 IRMAS dataset results
The SDR and ODG results for the IRMAS dataset are shown in Fig. 2. The reference B-PHADQ (oracle) shows the best results in both SDR and ODG, as expected. CP outperforms the proposed methods in terms of SDR. However, the consistent B-PHADQ performs no worse than by 1.3 dB across all word lengths. In terms of ODG, PHADQ proved to be superior, especially at a word length of 8 bps, where both consistent and inconsistent B-PHADQ exceeded CP by more than 0.3. Fig. 3 depicts the iteration number resulting in the best SDR and ODG reconstruction. At a word length of 8 bps, both B-PHADQ variants required no more than 60 algorithm iterations on average.
@@ 4.1 IRMAS dataset results
Fig. 3: Average number of iterations achieving the best SDR and ODG on the IRMAS dataset.
4.2 EBU SQAM dataset results
The results of the EBU SQAM dataset are presented Fig. 4 in the same way. There are some differences compared to the IRMAS dataset. The overall SDR is higher across all word lengths, whereas ODG shows lower values. This effect is due to the fact that the EBU SQAM dataset contains solo instruments, in which the (de)quantization error is hardly masked. In terms of SDR, CP again proves to be better than both consistent and inconsistent B-PHADQ. In terms of ODG, the B-PHADQ variants perform better than CP for most word lengths; however, the difference is not that pronounced than in the case of the IRMAS dataset.
Fig. 4: Average SDR and ODG on the EBU SQAM dataset
@@ 4.3 Computational considerations
For a 6-second-long excerpt, a single iteration of the consistent B-PHADQ variant takes about 0.12 second on a PC with AMD 3.80 GHz CPU and 32 GB RAM. Therefore, at a word length of 8 bps, optimal performance in terms of the ODG was reached in less than 7 seconds. A single iteration of the sparsity-based CP variant takes about 0.03 second, so a full reconstruction with 500 iterations requires approximately 15 seconds.
@@ 5 Conclusion
In this paper, we introduced a dequantization method that employs instantaneous frequency as a regularizer. The proposed approach demonstrated strong performance in objective evaluations, surpassing the reference method in terms of the perceptually-motivated ODG scale. Moreover, the method requires fewer algorithmic iterations than the sparsity-based CP algorithm, leading to reduced computational time. These findings highlight the potential of the proposed framework for efficient and high-quality restoration of quantized signals.

### v2

@@ Abstract
We present a dequantization method that employs a phase-aware regularizer, originally successfully applied in an audio inpainting problem. The method promotes a temporal continuity of sinusoidal components in time-frequency representation of the audio signal, and avoids energy loss artifacts commonly encountered with $\ell_{1}$-based regularization approaches. The proposed method is called the Phase-Aware Audio Dequantizer (PHADQ). The method are evaluated against the state-of-the-art using the SDR and PEMO-Q ODG objective metrics, and a subjective MUSHRA-like test.
@@ 1 Introduction
Quantization is a nonlinear distortion that arises from rounding during the A/D conversion or bit-depth reduction. This distortion appears as the quantization noise, which, at low bit depths (less than 16 bits per audio sample), results in unpleasant auditory artifacts [13]. Dequantization is the attempt to recover a signal close to its original form, based on the observed quantized samples. The problem is ill-conditioned; therefore, the reconstruction involves a form of regularization. For instance, is the application of sparsity [14] is one common approach. Autoregressive modeling is another modeling approach to dequantization [10, 12, 11].
In this paper, we present a dequantization method that employs a phase-aware regularizer introduced in [9], originally successfully applied in the context of the audio inpainting problem. The regularizer maintains the temporal continuity of sinusoidal components in the audio signal time-frequency representation and avoids the energy loss artifacts commonly encountered with $\ell_{1}$-based regularization approaches [9, 6]. The goal of this paper is to investigate whether the phase-aware regularization is beneficial also in the dequantization setup. The reconstruction quality and computational demands are compared to a baseline audio dequantization method [14].
@@ 2 Review of PHAIN (Phase-aware audio inpainter)
Problem (1) contains a penalty $\|DR_{\omega_{\mathbf{s}}}G_{g}{\mathbf{x}}\|_{1}$ weighted by a positive $\lambda$, and an indicator function $\imath_{\mathit{\Gamma}}({\mathbf{x}})$ taking value $\infty$ when ${\mathbf{x}}\notin\Gamma$, thus enforcing the estimation to retain the reliable parts of the signal corresponding to the set of feasible signals $\mathit{\Gamma}$}. The linear operators in the penalty function are:
  • The discrete Gabor transform (DGT) $G_{g}$ employing a window $g$.
@@
    where $a$ denotes the hop-size of the Gabor transform, and $n,m$ represent the time and frequency indices, respectively. The size of $\omega_{\mathbf{s}}$ is identical to the size of the spectrogram obtained by $G_{g}$.
@@
The convex $\ell_{1}$-norm denoted $\|\cdot\|_{1}$ quantifies time-directional variation of the phase-corrected time-frequency representation of $\mathbf{x}$}. When used as a penalty, it promotes sinusoidal components and, conversely, attenuates components whose phase cannot be adequately approximated as a linear function, such as noise or transients. As a consequence, PHAIN to a great extent solves the problem encountered in plain sparsity-based methods which is loosing signal energy inside the gap [6]. When $\omega_{\mathbf{s}}$ is fixed, the problem (1) is convex and solvable via the Chambolle–Pock algorithm [2, 3].
Based on the regime of treating the instantaneous frequency during the process of reconstruction, PHAIN can further be classified as follows:
  • B-PHAIN, which uses the instantaneous frequency $\omega_{\mathbf{y}}$ of the degraded signal, i.e. including unreliable parts.
  • U-PHAIN, which updates the instantaneous frequency according to the estimate of $\hat{{\mathbf{x}}}$ during iterations.
@@ 3 Method
We call the proposed method simply the Phase-Aware Audio Dequantizer (PHADQ). In the method, we employ the same penalty as in PHAIN, but for a different restoration task.
Let ${\mathbf{y}}^{\text{q}}\in\mathbb{R}^{L}$ be a quantized observation of an original audio signal with a quantization step $\Delta$. When employing uniform quantization, the quantization step size is fixed and given by $\Delta=2^{-w+1}$, where $w$ denotes the word length in bits per sample (bps). We recover the signal by solving the same convex problem as in (1), except that now no reliable parts of the signal are available. Therefore, in PHADQ, the set of feasible signals is characterized by:
@@
By choosing whether to update the instantaneous frequency estimate, we distinguish between the B-PHADQ and U-PHADQ, respectively.
@@ 3.1 B-PHADQ
Both the consistent and inconsistent restoration problems (i.e., (1) and (5), respectively) are solved by Algorithm 1, which constitutes a specific instance of the Chambolle–Pock algorithm, wherein ${\mathbf{x}}\in\mathbb{R}^{L}$, ${\mathbf{p}}\in\mathbb{R}^{L}$ is a primal variable, ${\mathbf{q}}\in\mathbb{C}^{P}$ is a dual variable and ${\mathbf{u}}\in\mathbb{R}^{L}$ is a temporary variable. Positive scalars $\tau$, $\sigma$ are the step sizes and $\rho$ is the relaxation factor [3]. The scalar $\lambda>0$ is the weighting factor for the penalty function, having also an effect on the reconstruction quality and the convergence speed. Before applying the algorithm, an initial estimate of the instantaneous frequency $\omega_{\mathbf{s}}$ has to be established, which corresponds to the following formula using the quantized observation [9]:
$\omega_{\mathbf{s}}[m,n]=\omega_{\mathbf{y}^{\text{q}}}[m,n]=-\text{Im}\Bigg[\cfrac{(G_{g^{\prime}}{\mathbf{y}^{\text{q}}})[m,n]}{(G_{g}{\mathbf{y}^{\text{q}}})[m,n]}\Bigg].$ (6)
Here, the Gabor transform is used twice, once with window function $g$ and the second time with $g^{\prime}$, the time derivative of $g$. $\text{Im}[\cdot]$ denotes the imaginary part.
@@
The algorithm utilizes proximal operators of the functions present in the minimization, which in the case of the $\ell_{1}$ norm correspond to the soft-thresholding, and in the case of the consistent indicator function it is the projection onto the set $\Gamma$. For the inconsistent case it is given by
@@
The asterisk in Algorithm 1 denotes the adjoint of a linear operator. Note that the term ${\mathbf{p}}^{(i+1)}$ is computed differently for the consistent and for the inconsistent variant. Finally, $\text{clip}_{\lambda}$ is defined as $\text{clip}_{\lambda}({\mathbf{x}})={\mathbf{x}}-\text{soft}_{\lambda}({\mathbf{x}})$.
@@ 3.2 U-PHADQ
As in U-PHAIN [9], an alternative approach to problems (1) and (5) is to regularly update the instantaneous frequency that is fixed in B-PHADQ in Algorithm 1. In U-PHADQ, such an update is performed after a user-prescribed number of iterations of Algorithm 1. The new instantaneous frequency is calculated for $\mathbf{p}$} rather than from $\mathbf{y}^{\text{q}}$, this way corresponding to the current state of reconstruction. U-PHADQ therefore contains the main loop of Algorithm 1 as its inner loop, the same way as in U-PHAIN [9].
@@ 4 Experiment and Results
Audio has been peak-normalized before any processing to make the most of the available dynamic range. The audio was quantized to seven different word lengths, ranging from 2 to 8 bps using the mid-riser uniform quantization. In the experiment, the Gabor transform with a Hann window of 8,192 samples in length, 75% overlap and 16,384 frequency channels has been used. All computation was performed in Matlab R2025a.
The parameters of Algorithm 1 were manually tuned. We used $\tau=1,\ \sigma=1$ and $\rho=1/3$. For better results, the balancing parameter $\lambda$ of the task needs to be set depending on the word length and variant; we set it as shown in Table 1. Indices c or i denote consistent or inconsistent variant. Note that $\lambda$ increases with decreasing bps, which is expected since the lower bps, the greater is the amount of error/noise. Also note that the inconsistent case requires lower $\lambda$ than its consistent counterpart, which makes sense, because in (5), the error is partially absorbed by the distance term, opposite to the hard constraint in (1).
Table 1: Values of $\lambda$ according to word length and method
@@
$\lambda_{\text{c}}$ $0.07$ $0.070$ $0.030$ $0.010$ $0.0010$ $0.0050$ $0.0002$
$\lambda_{\text{i}}$ $0.07$ $0.015$ $0.006$ $0.001$ $0.0008$ $0.0005$ $0.0002$
In the experiments, the performance of the methods was evaluated using the objective metric SDR, which expresses the signal-to-distortion ratio and measures the physical similarity of waveforms. We also employed the PEMO-Q objective difference grade (ODG) [4], which considers the perceptual characteristics of the audio signals. The ODG is measured on a scale from 4-4 to 0 (worst to best). To obtain subjective quality assessments, we employ the MUSHRA listening test [8].
@@
We show an example showing the evolution of SDR across iterations of PHADQ variants presented in Fig. 1. When comparing the U-PHADQ and B-PHADQ variants, it was found that updating the instantaneous frequency within the course of iterations of U-PHADQ does not lead to better results in terms of the SDR. For this reason, in the experiments we focus only on the B-PHADQ variant. In addition, to see the maximum achievable performance of PHADQ, there is the B-PHADQ (oracle) variant, based on the instantaneous frequency of the original signal, which cannot be accessed in practical scenarios. For the oracle variant the consistent B-PHADQ was used.
The results of the PHADQ method for both datasets show ODG and SDR after 60 iterations of the algorithm, which represents a balance between the average number of iterations required to reach optimal SDR and ODG values. We note that a higher number of algorithm iterations does not guarantee improved ODG values; the suplementary graphs are publicly available [^3] https://vojtechkovanda.github.io/PHADQ/. The CP results were obtained after 500 iterations of the algorithm, achieving optimal ODG values [14].
@@
4.1 Objective evaluation
The SDR and ODG results for the EBU SQAM and IRMAS dataset are shown in Fig. 2 and Fig. 3 respectively. As expected, the reference B-PHADQ (oracle) exhibits the best results in ODG in both datasets. CP outperforms the proposed methods in terms of SDR. However, in terms of ODG, the consistent B-PHADQ outperforms CP across almost all word lengths in both datasets. In the EBU SQAM dataset, all methods achieve higher SDR but lower ODG. This is due to the fact that EBU SQAM contains recordings of solo instruments, making the differences between recordings more perceptible. On the other hand, the IRMAS dataset contains recordings with multiple instruments, which masks the quantization noise more effectively, resulting in higher ODG.
Fig. 2: Average SDR and ODG on the EBU SQAM dataset
@@ 4.1 Objective evaluation
4.2 Subjective evaluation
The subjective test was conducted on the EBU SQAM database, comparing the CP, consistent B-PHADQ, and inconsistent B-PHADQ methods. Only 6-bit and 7-bit word lengths were used in the test, as these are the highest bit depths at which quantization distortion is still relatively clearly perceptible. The test results are shown in Fig. 4 and confirm the objective evaluation.
Fig. 4: MUSHRA results for 6 and 7 bps quantization on EBU SQAM dataset
@@ 4.3 Computational considerations
For a 6-second-long excerpt, a single iteration of the consistent B-PHADQ variant takes about 0.07 second on a PC with AMD 3.80 GHz CPU and 32 GB RAM. Therefore a full, 60 iteration long reconstruction is obtained in 4.2 seconds. A single iteration of the sparsity-based CP variant takes about 0.03 second, thus a full reconstruction with 500 iterations requires approximately 15 seconds.
@@ 5 Conclusion
In this paper, we introduced a dequantization method that employs instantaneous frequency as a regularizer. The proposed approach demonstrated strong performance in objective evaluations. Moreover, the method requires fewer algorithmic iterations to converge than the sparsity-based CP algorithm, leading to reduced computational time. These findings highlight the potential of the proposed framework for efficient and high-quality restoration of quantized signals.